OpenStudy (anonymous):

11 to what power gives me 1 over the square root of 11

OpenStudy (anonymous):

Are you sure this is worded correctly?

OpenStudy (espex):

Is the question, \(11^x=\frac{1}{\sqrt{11}}\)?

OpenStudy (espex):

If so, recall that a root can be expressed as a fraction, and that powers in the denominator can be expressed as a negative of that power.